Circular No. 8757

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 T1 (LEVY) David H. Levy, Tucson, AZ, reports his visual discovery with his 0.41-

m f/5 reflector of a diffuse comet that is close to Saturn in the sky; the discovery position tabulated below is from CCD astrometry obtained with a 0.36-m reflector, while the magnitude is visual. Following posting on the ‘NEO Confirmation Page’, other observers have confirmed the object as a comet, including K. Sárneczky (Szeged, Hungary, 0.60-m Schmidt telescope), who reports that his CCD frames taken on Oct. 3.13–3.14 UT show a strong central condensation of magnitude 15 and a coma of diameter 1′. P. Birtwhistle (Great Shefford, Berkshire, U.K., 0.40-m f/6 Schmidt-Cassegrain telescope) writes that his CCD images taken on Oct. 3.2 show a circular coma of diameter 4′.5 with a concentrated but non-stellar center and a thin, straight tail 14′ long in p.a. 295^{o} (10′′ wide as it leaves the coma, fanning out to 2′ wide as it fades into the sky background). R. Miles (Stourton Caundle, Dorset, U.K., 0.06-m refractor) communicates that his CCD frames taken on Oct. 3.2 show a total V magnitude of 9.5 in a 3′ aperture. E. Guido and G. Sostero (observing remotely with a 0.25-m f/3.4 reflector near Mayhill, NM) write that their CCD images from Oct. 3.48 show a coma diameter of 3′ and a tail 2′ long toward p.a. 297^{o}. A. Hale, Cloudcroft, NM (0.41-m reflector) reports that his visual observation on Oct. 3.47 yields a coma diameter of 2′.5 and total magnitude 9.8. J. E. McGaha (Tucson, AZ, 0.36-m f/10 Schmidt-Cassegrain reflector + CCD) reports that six stacked 60-s frames taken on Oct. 3.5 show a bright, round, inner coma of diameter 34′′ with an outer coma of diameter 226′′, elongated toward a 30′′-wide tail that extends 600′′ in p.a. 287^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
Oct. 2.50061  9^{h}37^{m}29.^{s}47  +15^{o}52′43′′.1  10.5

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-T21.

T = 2006 Oct. 9.226 TT  ω = 181.^{o}745
Ω = 284.416 }2000.0
q = 1.07182 AU  i = 19.117
2006 October 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green